# Nonlinear Impulse Pattern Formulation dynamical social and political prediction algorithm for city planning and public participation

### v1

@@ III.4.1 Complex delay interactions
In the methods, we presented a more complex version of the mediating interaction. Eq. 7 describes the change of $\beta^{s,i}$ as caused by a previous state of stakeholder $i$ on a previous state of stakeholder $s$, not on the present state of stakeholder $s$ like in Eq. 6. This adds a memory to the stakeholders as they remember their previous states and take this into account while judging the impact of a coupled stakeholder.
@@
Figure 17: Trained SOM of mediated interaction but with more complex delay term using Eq. 7: a) and b) show the same u-matrix, dots represent the best-fit IPF cases onto the map. a) Colors are chosen according to initial conditions. Red dots: I1 and I2, white dots: I3 and I4. b) Colors represent coupling strength $V$, with darker colors representing weaker, and lighter colors representing stronger coupling.
@@
Still, the system is more unstable. Fig. 18 shows that using Eq. 7, the chaotic transients can be significantly extended, and the IPF must not converge anymore.
Figure 18: Time series of $g^{s}$, for each of the s=6 stakeholders of the mediated interaction Version II of Varying terms III but here with complex delay of Eq. 7 with initial condition I3, $\alpha^{s}\approx\{0.57,0.94,0.50,0.71,0.94,0.97\}$, $V\approx 0.008$, and $C_{dt}=8$. The case is mapped to the upper left corner of the SOM shown in Figure 17.
@@ IV Conclusions
The proposed algorithm is able to model all typical kinds of city planning, social, or political time developments of interacting and self-adapting stakeholders. These include conversion to a common state, conversion to different states, bifurcations of many kinds, as well as chaotic behaviour. The reaction to an external impact is also realistic with adaptation, convergence, or breakdown.
@@
The model is flexible in terms of using different varying Terms in different variations. This allows flexible modification of the model according to real-world situations. Also, by adding or omitting terms, certain kinds of interactions or self-adaptations might be suggested by the system, therefore making aware of critical stakeholder interaction processes needing more care during a planning process.

### v2

@@ III.4.1 Complex delay interactions
In the methods Section, we presented a more complex version of the adaptive interaction. Eq. 10 describes the change of $\beta^{s,i}$ as caused by a previous state of stakeholder $i$ on a previous state of stakeholder $s$, not on the present state of stakeholder $s$ like in Eq. 9. This adds a memory to the stakeholders as they remember their previous states and take this into account while judging the impact of a coupled stakeholder.
@@
Figure 17: Trained SOM of adaptive interaction but with more complex delay term using Eq. 10: a) and b) show the same u-matrix, dots represent the best-fit IPF cases onto the map. a) Colors are chosen according to initial conditions. Red dots: I1 and I2, white dots: I3 and I4. b) Colors represent coupling strength $V$, with darker colors representing weaker, and lighter colors representing stronger coupling.
@@
Still, the system is more unstable. Fig. 18 shows that using Eq. 10, the chaotic transients can be significantly extended, and the IPF may not converge anymore.
Figure 18: Time series of $g^{s}$, for each of the N=6 stakeholders of the adaptive interaction Version II of Varying terms III but here with complex delay of Eq. 10 with initial condition I3, $\alpha^{s}\approx\{0.57,0.94,0.50,0.71,0.94,0.97\}$, $V\approx 0.008$, and $C_{dt}=8$. The case is mapped to the upper left corner of the SOM shown in Figure 17.
@@ IV Conclusions
The proposed algorithm is able to model all typical kinds of city planning, social, or political time developments of interacting and self-adapting stakeholders. These include conversion to a common fixpoint, conversion to different states, bifurcations or oscillations of many kinds, as well as chaotic behaviour. The reaction to an external impact is also realistic with adaptation, convergence, or breakdown.
@@
The model is flexible in terms of using different varying terms in different versions. This allows flexible modification of the model according to real-world situations. Also, by adding or omitting terms, certain kinds of interactions or self-adaptations might be suggested by the system, therefore making aware of critical stakeholder interaction processes needing more care during a planning process.